You are here: Foswiki>PartnerObserving Web>ItalyQueue2013>ObsInaf20130928 (29 Sep 2013, EleanoraSani)Edit Attach

INAF queue observing, 2013 Sep 27-28

Observer: E. Sani, F. Cusano
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: D. Thompson (on call)
Instrument: LUCI

Summary

MODS1 is currently available in red or blue grating only.
At the sunset partially cloudy and temperature close to due point.
When the TO opens the dome we ave clear sky and low wind.
At the beginning of scientific integtrartion the average seeing on the GCS is 1.7" with 90% humidity.
During the night we had 2 issue tracks, one for stuck rotaror and the over to shrink condensation from LUCI.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Cresci_8B Hip107734 210_zJHK, H, K   bad seeing conditions K CW=2.17
Cresci_8B DSF2237b 210_zJHK, K 1h bad seeing at the beginning. Totatal oh 5h of integration reached, but some frames to be repeated. K CW=2.17
Cresci_8B DSF2237b 210_zJHK, H 2h variable seeing, but mostly within or close to the required 1"
Cresci_8B Hip18874 210_ZJHK, H, K   K CW=2.17
Galametz_31A Hip11809 200_H+K, zJspec   condensation. Loose spectra for slits from #11 to #32
Galametz_31A CANDELS_UDS18 200_H+K, zJspec 10m condensation. Loose spectra for slits from #11 to #32

Details

02:54 The TO opens the dome

03.17 We slew to standard star for Cresci_28B
  • preset cancelled due to technical operations on LUCI2.

03.26 execureLUCI tell2B_mos.acq 0001-0003

03.37 execureLUCI tellB_mos_210_K1.spec
  • there are problems in stabilizing the 210_zJHK grating central wavelength. The SA has to intervene through the engineering interface.

04.01 tellB_mos_210_H2.spec 0004-0007
  • the problem is not solved, we try to first run the Hband script

04.04 execureLUCI tellB_mos_210_K1.spec
  • we tried again with the CW setted by the PI, but it failed. We thus change the CW from 2.15 mu to 2.17 mu

04.15 execureLUCI tellB_mos_210_K1.spec 0008-0011

04.19 We slew to the target DSF2237b
  • execureLUCI campoB_mos.acq
  • with a guiding star R=13.4 we are not able to guide due to bad seeing 2.0 average on the GCS

04:34 We ask the TO to slew to a bright star to try to better collimate
  • In the meanwhile humidity is reaching 90% and the temperature is approacing the dewpoint.

04.40 execureLUCI campoB_mos.acq 0012-0015
  • the average seeing on the GCS is 1.5-2.0

04:58 execureLUCI campoB_mos_210_K_5.spec 0016-0027
  • with CW 2.17 mu
  • decreasing humidity
  • the average seeing on the GCS at the beginning of integration is 1.5. During the integration it decreases beeing 1.1 at the 4th offset.
  • the average seeing on the GCS at the end of integration is 1.2

06:05 execureLUCI campoB_mos_210_H_1.spec 0028-0031
  • only 4 offsets acquired because of a preset cancelled. The rotator is stuck

06:33 execureLUCI campoB_mos.acq 0032-0035

06.47 execureLUCI campoB_mos_210_H_1.spec 0036-0043
  • seeing 0.9"-1.2"

07:34 execureLUCI campoB_mos_210_H_2.spec 0044-0055
  • Thin cirrus passing
  • Humidity is increasing again
  • High seeing variability: the average value on the GCS range is 0.8"-1.7"

08:41 execureLUCI campoB_mos_210_H_3.spec 0056-0057
  • WARNING level reached for sensor shield temperature
  • Airmass limit reached after 2 exposures.

08:53 slewing to standard star HIP116324
  • executeLUCI tellB3_mos.acq
  • we are not guiding for bad seeing
  • Collimating on a bright star
  • 0058 test image

09:16 executeLUCI tellB3_mos.acq
  • seeing 2-3, conditions are too bad to collimate. Humidity is 92% thus within the critic limit, but we cannot start the integration since the telescope is still commited, moreover we cannot close the dome because the temperature is still higher than the dewpoint

09.23 executeLUCI tellB3_mos.acq 0059
  • we have to exit the script because J. Morris has to reach the LUCI platform

09:30 executeLUCI tellB3_mos.acq 0060-0062
  • not able to collimate
  • seeing 1.5

09:45 the TO closes the dome, condensation on the LUCI camera due to persistent (more than 1h) humidity at 92%

10:12 Telescope in rest position. J. Morris is on the LUCI platform to increase air pressure and dry condensation.
  • We take 4s images to monitor the status of condensation 0063-0065

10.23 The TO opens the dome. We continue to monitor LUCI
  • We take 4s images to monitor LUCI 0066, there is still condensation on the camera

10:36 slew to standard of Galametz_31
  • executeLUCI CANDELS_UDS18_A1_tell_mos.acq 0067-0071
  • we loosed time to acquire the standard because it was folling in the condensation area and was not possible to identify it.

11:09 executeLUCI CANDELS_UDS18_A1_tell_mos_zJspec.spec 0072-0074
  • the spectreum on the first slit (#16) falls in the condensation shadow area. Slit #16 is a central slit. The reddest and bluest spectra are acquired. Data can be calibrated

11:11 executeLUCI CANDELS_UDS18_A1_mos.acq 0075-0078

11:26 CANDELS_UDS18_A1_mos_zJspec_2.spec 0079-0080
  • due to condensation we are loosing spectra from slit #11 to the #32 *airmass 1.55. Average seeing on the GCS: 1.5

11:37 slew to standard CresciB Hip18871
  • executeLUCI tell2B_mos.acq 0081-0082
  • seeing 1.6

11:48 executeLUCI tellB_mos_210_K1.spec 0083-0086

11:55 executeLUCI tellB_mos_210_H.spec 0087-0090

12:02 the TO closes the dome

12:28 starting calibrations for CresciB K&H bands
  • we have to repeat the calibration for Cresci of night 20130925 because the central wavelegth is set to a wrong value
  • executeLUCI campoB_flat_K.cal 0091-0100
  • executeLUCI campoB_flat_K_2.17.cal 0101-0110
  • executeLUCI campoB_flat_H.cal 0111-0120
  • executeLUCI campoB_arc_K.cal 0121-0124
  • executeLUCI campoB_arc_K_2.17.cal 0125-0128
  • executeLUCI campoB_arc_H.cal 0129-0132
  • images at 4s to check the condensation 0133-0134

-- FeliceCusano - 28 Sep 2013
Topic revision: r6 - 29 Sep 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback